OpenStudy (anonymous):

a landscaper charges $28 per hour of labor and $105for plants and materials to do small planting job . if the total cost is $168 how long did the job take show work

OpenStudy (anonymous):

let's call t= time to do the job so, we know that 105 +28t = 168 can you solve it?

OpenStudy (anonymous):

28t=168-105 28t=63 t=63/28 or it will take 2 hours & 15 min (1/4 of the hour)